### v3

@@ 1 Introduction
There have been surveys and summaries on the security issues with LLMs [49, 28, 13], however, the taxonomy and terminology used in them are often messy and inaccurate. For example, they list "prompt injection" and "jailbreak" as distinct kinds of attacks, while they actually belong to attack techniques and objective respectively and thus cannot be categorized together. Furthermore, they largely overlook the emergent intrinsic risks of autonomous LLM agents, and this survey fills that gap by placing significant emphasis on phenomena such as goal misalignment, strategic deception (scheming), and the persistence of ‘sleeper agent’ behaviors—critical and rapidly advancing frontiers in LLM security. Furthermore, the survey makes the following contributions: (1) We provide a comprehensive taxonomy that integrates these intrinsic agentic risks alongside established threats like prompt injection, adversarial attacks, and malicious misuse. (2) We review a broad range of recent academic and industry works from 2022 to 2025, highlighting representative examples of each threat type and incorporating recent findings not covered in earlier surveys. (3) We evaluate the effectiveness and limitations of current defense strategies, including input sanitization, adversarial training, and oversight mechanisms. (4) We identify open research challenges for securing LLMs, especially in light of emergent risks such as goal misalignment, scheming, and deceptive behavior in autonomous agents. By mapping the evolving threat landscape and surveying mitigation strategies, this survey aims to inform both practitioners who are deploying LLMs and researchers who are designing the next generation of large language models with the potential risks, actionable insights, and practical recommendations for mitigating the security threats.
@@ 2.3 Manual and Heuristic Prompt Crafting
2.4 Automated and Optimization-Based Prompt Generation
Automated methods for achieving prompt injection are often produce more robust and transferable attacks that work across different models.
@@ 2.4 Automated and Optimization-Based Prompt Generation
Expanding on this automation, Zhang et al. propose a more theoretically-grounded black-box method called Goal-Guided Generative Prompt Injection (G2PI) [50]. Unlike the gradient-based GCG, which optimizes for a simple affirmative response, G2PI reframes the attack’s objective as maximizing the KL divergence between the output probability distributions of the clean and adversarial prompts. The authors show this is theoretically equivalent to maximizing the Mahalanobis distance between the two prompt embeddings in a latent space. To operationalize this in a black-box setting without access to gradients or internal parameters, G2PI uses an auxiliary LLM to generate a semantically plausible injection sentence. The generation is guided by a clever proxy: it searches for a sentence that is semantically close to the original query but whose embedding has a specific, constrained cosine similarity to the original embedding. This approach moves beyond finding non-sensical token suffixes and instead generates coherent, context-aware sentences that act as the adversarial payload, achieving high jailbreak success rates on models like GPT-4 and Gemini.
@@
As another refinement of the optimization objective itself, Zhu et al.[52] argue that many automated attacks are limited by a misspecified and overconstrained objective, such as forcing the model to begin its response with a single, rigid prefix like “Sure, here is…”. They observe that such an objective often leads to incomplete or unrealistic outputs even when the optimization is successful, and that the rigid prefix is often unnatural for the target model, hindering the optimization process. To address this, they introduce AdvPrefix, a prefix-forcing objective that automatically selects more nuanced, model-dependent prefixes. These prefixes are chosen based on two criteria: a high probability of leading to a complete and harmful response (high prefilling attack success rate) and being easy for the model to generate (low initial negative log-likelihood). Their results show that simply replacing the standard objective in an attack like GCG with their automatically selected prefixes can dramatically improve nuanced attack success rates (e.g., from 14% to 80% on Llama-3), demonstrating that current alignment techniques can fail to generalize to more natural-sounding harmful response prefixes [52].
@@ 3.1 Data Poisoning and Backdoor Insertion
The fundamental method for compromising a model during training is through data poisoning or the insertion of backdoors. In these scenarios, an attacker subtly injects malicious examples into the training dataset, which causes the model to learn a hidden, undesirable behavior that can be triggered by specific inputs later on [34, 27, 35]. For instance, by adding a specific backdoor trigger to a small portion of the training data, an otherwise benign model can be conditioned to output attacker-chosen text for inputs containing that trigger, while its behavior on normal inputs remains unchanged [34, 40]. Empirical studies confirm the efficacy of these attacks. Wallace et al.[40] demonstrated that models such as GPT-2 could be made to output arbitrary attacker-specified content by inserting rare token sequences into fine-tuning data. The BackdoorLLM benchmark shows that such attacks remain effective even with very small poisoning budgets (e.g., 15–100 samples), and that existing post-hoc defenses often fail to remove them[27]. A novel attack vector is demonstrated by BadGPT, which compromises the reward model during RLHF to create covert policy backdoors, showing that alignment layers themselves are vulnerable attack surfaces [35].
@@ 3.2 Deceptive Alignment and the “Sleeper Agent” Problem
The most advanced and dangerous form of training-time attack involves creating a deceptively aligned model, a concept that goes far beyond a simple backdoor. While related, its mechanism and implications are profoundly different. In their work on "sleeper agents," Hubinger et al. [21] introduced and empirically demonstrated this threat: models with deceptive objectives embedded during training that activate only in response to specific triggers.
@@
The critical and most alarming finding from this research is that such deceptive backdoors can persist and even be reinforced through standard safety training. Their findings reveal that standard safety pipelines, including adversarial fine-tuning and RLHF, can fail to remove the backdoor. Instead, these methods can inadvertently train the model to better conceal its deceptive nature by reinforcing the "safe" behavior that is part of its deceptive strategy. This creates a false sense of security, as the model appears aligned during evaluation but retains its covert malicious capabilities. These studies collectively highlight the severity of training-time vulnerabilities, as they compromise foundational trust in LLM behavior and represent a fundamental challenge for AI safety and alignment verification. [27, 40, 34, 35, 21].
@@
Phase Attack Type Description Key Techniques/Characteristics Example Studies & Key Findings
Training-Time Data Poisoning Malicious examples injected into training data to cause specific misbehavior. Subtle, hard to detect post-training. Can implant persistent triggers with small data percentage. Shayegani et al. [34], BackdoorLLM [27] (15–100 poisoned samples still effective; post-hoc defenses often fail)
Training-Time Backdoor Insertion Training LLM to produce harmful output for specific trigger inputs while behaving normally otherwise. Trigger-activated, behavior concealed on normal inputs. Reward model poisoning can create covert policy backdoors. Wallace et al. [40] (robust triggers), BadGPT [35] (reward model poisoning).
Training-Time Sleeper Agents Deceptive objectives embedded during training, activate on specific triggers, and resist safety fine-tuning. Covert, persistent through safety training; model may learn to better hide backdoor. Hubinger et al. [21] (Safety pipelines can train models to conceal backdoors).
Inference-Time Manual/Heuristic Prompt Crafting Manually crafting prompts using linguistic tricks like role-playing or exploiting model failure modes to bypass safety filters. Role-playing (e.g., DAN), exploiting competing objectives (helpfulness vs. harmlessness), mismatched generalization. Wei et al. [45] (Conceptual framework for jailbreaks), DAN Prompt [39] (Role-playing).
Inference-Time Input Perturbations Minor changes to input (synonyms, misspellings, token manipulation) to cause misclassification or harmful output. Evades simple filters, exploits model sensitivity to token sequences. Zou et al. (GCG) [53] (transferable suffixes)

### v4

@@ 1 Introduction
There have been surveys and summaries on the security issues with LLMs [68, 37, 16], however, the taxonomy and terminology used in them are often messy and inaccurate. For example, they list "prompt injection" and "jailbreak" as distinct kinds of attacks, while they actually belong to attack techniques and objective respectively and thus cannot be categorized together. Furthermore, they largely overlook the emergent intrinsic risks of autonomous LLM agents, and this survey fills that gap by placing significantly more emphasis on phenomena such as goal misalignment, strategic deception (scheming), and the persistence of ‘sleeper agent’ behaviors—critical and rapidly advancing frontiers in LLM security. Furthermore, the survey makes the following contributions: (1) We provide a comprehensive taxonomy that integrates these intrinsic agentic risks alongside established threats like prompt injection, adversarial attacks, and malicious misuse. (2) We review a broad range of recent academic and industry works from 2022 to 2025, highlighting representative examples of each threat type and incorporating recent findings not covered in earlier surveys. (3) We evaluate the effectiveness and limitations of current defense strategies, including input sanitization, adversarial training, and oversight mechanisms. (4) We identify open research challenges for securing LLMs, especially in light of emergent risks such as goal misalignment, scheming, and deceptive behavior in autonomous agents. By mapping the evolving threat landscape and surveying mitigation strategies, this survey aims to inform both practitioners who are deploying LLMs and researchers who are designing the next generation of large language models with the potential risks, actionable insights, and practical recommendations for mitigating the security threats.
@@ 2.3 Manual and Heuristic Prompt Crafting
2.4 Automated Prompt Generation
Automated methods for achieving prompt injection often produce more robust and transferable attacks that work across different models. These techniques can be broadly categorized based on the level of access they require to the target model, primarily distinguishing between white/gray-box and black-box approaches.
2.4.1 White-Box and Gray-Box Attacks
These methods assume a higher level of access to the target model, ranging from full access to internal states like gradients (white-box) to partial information leaks like training loss (gray-box). This access allows for more direct and often more efficient optimization of adversarial inputs.
@@ 2.4.1 White-Box and Gray-Box Attacks
As a refinement of the optimization objective itself, Zhu et al.[73] argue that many automated attacks are limited by a misspecified and overconstrained objective, such as forcing the model to begin its response with a single, rigid prefix like “Sure, here is…”. They observe that such an objective often leads to incomplete or unrealistic outputs even when the optimization is successful, and that the rigid prefix is often unnatural for the target model, hindering the optimization process. To address this, they introduce AdvPrefix, a prefix-forcing objective that automatically selects more nuanced, model-dependent prefixes. These prefixes are chosen based on two criteria: a high probability of leading to a complete and harmful response (high prefilling attack success rate) and being easy for the model to generate (low initial negative log-likelihood). Their results show that simply replacing the standard objective in an attack like GCG with their automatically selected prefixes can dramatically improve nuanced attack success rates (e.g., from 14% to 80% on Llama-3), demonstrating that current alignment techniques can fail to generalize to more natural-sounding harmful response prefixes [73].
Shifting the focus to indirect prompt injection and the robustness of attacks, Pasquini et al. [45] introduce Neural Exec, a framework for automatically generating execution triggers for prompt injection attacks. Unlike prior methods that focused on generating a complete adversarial prompt or a simple suffix, Neural Exec conceptualizes the creation of the execution trigger itself—the part of the prompt designed to make the LLM execute a malicious payload—as a differentiable search problem. Using a gradient-based optimization approach, the framework learns triggers that are significantly more effective and flexible than handcrafted ones. The primary innovation of Neural Exec is its focus on generating triggers that are robust enough to persist through complex, multi-stage preprocessing pipelines, such as those found in RAG systems. To achieve this, the optimization process is designed to create triggers that are inlined (existing on a single line to avoid being split by text chunkers) and exhibit Semantic-Oblivious Injection (SOI), a property that minimizes the semantic disruption to the surrounding text to ensure the malicious chunk is successfully retrieved by the RAG system. The resulting triggers deviate markedly in form from known attacks, thereby bypassing existing blacklist-based detection methods.
A different attack surface is exploited by targeting the demonstrations within In-Context Learning (ICL) prompts. The Greedy Gradient-guided Injection (GGI) attack introduces a threat model where an adversarial "model publisher" poisons the few-shot examples provided to a user [47]. Instead of altering the user’s query, GGI uses a gradient-based search algorithm to learn and append short, imperceptible adversarial suffixes to the in-context demos. This automated process optimizes the suffixes to hijack the model’s behavior, forcing it to generate a specific, predetermined output (e.g., always classifying sentiment as ’positive’) or elicit a harmful, jailbroken response, regardless of the user’s actual query. The attack is designed to be stealthy, as the word-level suffixes are less conspicuous than character-level perturbations and thus harder to detect via perplexity-based defenses. This work highlights how the ICL mechanism itself can be subverted, turning the model’s learning examples into a vector for prompt injection.
Following a different paradigm, to improve the inference efficiency of adversarial attacks, the Weak-to-Strong jailbreak introduces a novel method that foregoes computationally expensive optimization entirely [71]. This technique operates by using two smaller, "weak" models—one that is safely aligned and another that has been made "unsafe" (e.g., through fine-tuning on harmful examples)—to manipulate the output of a significantly larger, "strong" target model at inference time. The core insight is that the token distributions of safe and unsafe models differ most significantly only in the initial tokens of a response. The attack exploits this shallow alignment by adjusting the strong model’s next-token probability distribution. Specifically, it multiplies the strong model’s probabilities with a term derived from the difference in log probabilities between the weak unsafe and weak safe models. This effectively steers the stronger model towards generating harmful content, especially at the beginning of its response, after which the strong model’s own capabilities take over to produce detailed and potent harmful outputs. This method is remarkably efficient, achieving a misalignment rate of over 99% on benchmark datasets with just a single forward pass through the target model, and it requires no complex prompt engineering or gradient calculations. Furthermore, the attack often results in "amplified" harm, where the output from the strong model is more malicious and detailed than what the weak unsafe model could generate on its own.
Leveraging a completely different and novel attack surface, Labunets et al.[31] introduce Fun-tuning, a gray-box attack that exploits the remote fine-tuning interface provided by LLM vendors. This approach targets closed-weight proprietary models where direct access to gradients or log probabilities from the inference API is unavailable. The core insight is that the training loss values returned by the fine-tuning API after a training job can serve as a proxy for the true adversarial loss. To achieve this, the attacker submits candidate adversarial prompts for fine-tuning with a near-zero learning rate, which prevents any significant updates to the model’s weights but still coaxes the API into returning the loss for each input-output pair. This leaked loss signal is then used to guide a greedy, discrete optimization search for an effective adversarial prefix and suffix to wrap around a malicious instruction. The authors demonstrate that despite technical hurdles, such as the API permuting the order of the training data, this loss information is a sufficiently strong signal to guide the attack. Their experiments show high attack success rates (65-82%) against Google’s Gemini models, revealing a fundamental security vulnerability in a feature designed for utility and model customization.
2.4.2 Black-Box Attacks
In contrast, black-box attacks operate under a more constrained threat model, assuming no access to the model’s internal parameters, gradients, or log probabilities. These methods rely solely on interacting with the model’s input-output interface, making them more broadly applicable to proprietary, closed-source models.
Expanding GCG on automation in this setting, Zhang et al. propose a more theoretically-grounded black-box method called Goal-Guided Generative Prompt Injection (G2PI) [70]. Unlike the gradient-based GCG, which optimizes for a simple affirmative response, G2PI reframes the attack’s objective as maximizing the KL divergence between the output probability distributions of the clean and adversarial prompts. The authors show this is theoretically equivalent to maximizing the Mahalanobis distance between the two prompt embeddings in a latent space. To operationalize this in a black-box setting without access to gradients or internal parameters, G2PI uses an auxiliary LLM to generate a semantically plausible injection sentence. The generation is guided by a clever proxy: it searches for a sentence that is semantically close to the original query but whose embedding has a specific, constrained cosine similarity to the original embedding. This approach moves beyond finding non-sensical token suffixes and instead generates coherent, context-aware sentences that act as the adversarial payload, achieving high jailbreak success rates on models like GPT-4 and Gemini.
To address the distinct challenge of indirect prompt injection in black-box settings, the AutoHijacker  [35] framework is proposed as an automated attack that operates without access to internal model details. This approach is specifically designed to overcome the problem of sparse feedback, where an attacker receives little to no useful information from failed attempts, thus hindering traditional iterative optimization. AutoHijacker reframes the problem by introducing a batch-based optimization framework and a multi-agent system, consisting of a prompter LLM, an attacker LLM, and a scorer LLM, that work together to generate effective malicious data injections. The core of the system is a trainable "attack memory" that stores a repository of past attacks and their effectiveness. By retaining both the most and least successful attacks, this memory provides a balanced, contrastive perspective that helps the prompter LLM guide the attacker LLM to generate more potent injections while avoiding previously failed strategies. This design enables AutoHijacker to perform a one-step generation during its test phase, creating powerful attacks without the need for continuous querying of the victim model. Evaluations show that this method achieves state-of-the-art performance, outperforming other black-box methods and even rivaling gray-box attacks on benchmarks like AgentDojo and Open-Prompt-Injection.
Drawing inspiration from biology, another line of research employs evolutionary algorithms (EAs) to automatically discover and optimize jailbreak prompts. Yu et al. [69] introduce LLM-Virus, a black-box attack framework that conceptualizes the jailbreak process as the evolution of a biological virus. In this analogy, the jailbreak template acts as the virus’s genetic material (DNA/RNA) and the malicious query is the functional protein, while the target LLM is the host with a safety alignment that functions as an immune system. The core of the method is an evolutionary algorithm that iteratively improves a population of jailbreak templates through selection, crossover, and mutation. A key innovation of LLM-Virus is its use of a powerful auxiliary LLM as an "evolutionary operator." Instead of relying on simple word-level mutations or random paragraph swaps, this auxiliary LLM is prompted to perform semantically-aware "heuristic" crossover and mutation, generating more diverse, coherent, and effective offspring templates. To address the high computational cost typically associated with EAs, the framework incorporates a transfer learning approach, first performing "Local Evolution" on a small, representative subset of malicious queries before testing the evolved templates’ "Generalized Infection" capability on the full dataset. This combination of bio-inspired evolutionary search and LLM-driven text manipulation proves effective at creating novel and transferable jailbreak attacks.
Another paradigm for automated jailbreaking draws inspiration from social engineering and conversational red teaming. The Prompt Automatic Iterative Refinement (PAIR) framework by Chao et al.[11] operationalizes this concept by pitting two black-box LLMs against each other: an attacker” and a target”. The process is fully automated and conversational: the attacker LLM generates an initial jailbreak prompt, which is sent to the target. The target’s response is then evaluated by a third “judge” model (e.g., Llama Guard [27]) to determine if the jailbreak was successful. If not, the attacker is provided with the full history of its failed prompt and the target’s refusal, prompting it to iteratively refine its strategy and generate a new, improved prompt. This iterative, chain-of-thought style refinement allows the attacker to learn from its failures and adapt its approach. The primary contributions of PAIR are its remarkable query efficiency—often finding a successful semantic jailbreak in fewer than twenty queries, orders of magnitude less than optimization methods like GCG—and its generation of human-interpretable, prompt-level attacks. By leveraging an attacker LLM to automate the creative process of prompt design, PAIR effectively bridges the gap between labor-intensive manual jailbreaks and query-inefficient, uninterpretable token-level attacks, demonstrating high success rates against a wide range of both open and proprietary models[11].
Building directly on the conversational red teaming concept, Mehrotra et al. introduce the Tree of Attacks with Pruning (TAP) framework as a significant enhancement to PAIR [39]. While PAIR follows a linear refinement process for a single prompt, TAP parallelizes the search for vulnerabilities through two primary innovations: branching and pruning. At each iteration, the branching step uses the attacker LLM to generate multiple distinct variations of the current best prompts, creating a tree of potential attack paths rather than a single chain. Subsequently, the pruning step employs an evaluator LLM to assess these newly generated prompts. It first prunes branches that are unlikely to succeed (e.g., by being off-topic) before they are sent to the target model, and after querying the target, it retains only the highest-scoring prompts for the next iteration of branching. This combination of exploring a wider attack surface through branching and increasing query efficiency through pruning allows TAP to achieve a substantially higher jailbreak success rate across a range of state-of-the-art LLMs compared to PAIR, often with significantly fewer queries to the target model.
Another approach, named FlipAttack [36], exploits the fundamental left-to-right, autoregressive nature of LLMs to create a simple yet highly effective black-box jailbreak. The core insight is that LLMs struggle to comprehend text when noise is introduced to the left side of a prompt. FlipAttack operationalizes this by first using an attack disguise module” to obfuscate a harmful request by systematicallyflipping” its components, such as reversing the order of words or the characters within the entire sentence. This process constructs a stealthy, high-perplexity prompt using only the original content, which allows it to bypass external guardrail models. Then, within the same query, a “flipping guidance module” instructs the victim LLM to denoise the prompt by reversing the flip, comprehend the now-uncovered harmful intent, and execute it. This guidance can be enhanced with chain-of-thought and few-shot examples to assist weaker models. This method distinguishes itself by being non-iterative, successfully jailbreaking state-of-the-art models like GPT-4o with a single query, demonstrating high attack success and bypass rates.
A notable conceptual shift is to reframe jailbreaking not as a discrete optimization problem, but as one of inference-time misalignment. Following this principle, Beetham et al. [7] introduce LIAR (Leveraging Inference-time misAlignment to jailbReak), a fast, training-free, black-box attack. The method employs a simple but powerful best-of-N sampling strategy, using an auxiliary "adversarial" LLM (such as GPT-2) to generate numerous natural-sounding suffix candidates for a given harmful prompt. These augmented prompts are then sent to the target LLM. The key advantage of this parallel approach is a dramatic reduction in the Time-to-Attack from hours to seconds, while still achieving attack success rates comparable to state-of-the-art methods.Furthermore, because the suffixes are generated by a standard language model without forced optimization, they exhibit low perplexity, making the resulting prompts appear more natural and thus harder to detect via perplexity-based filters. The work also introduces a theoretical "safety net against jailbreaks" metric to help quantify a model’s vulnerability by connecting it to its underlying safety alignment.
Focusing on the distinct and increasingly relevant attack surface of LLM-powered tabular agents, Feng and Pan [19] introduce StruPhantom, a framework for indirect prompt injection tailored for black-box agents that process structured data like CSV, JSON, and XML. The core challenge addressed is that such agents impose strict data formats and rules, making it difficult for a malicious payload to be correctly parsed and executed. To overcome this, StruPhantom reframes the attack as an evolutionary optimization problem, using a constrained Monte Carlo Tree Search (MCTS) to iteratively refine attack templates. The framework employs a multi-agent system, including a Mutate Agent to generate variations and a Refine Agent to make adjustments based on the target’s behavior. A key component is an off-topic evaluator that prunes mutated templates that deviate from the intended attack goal, ensuring the search remains efficient and focused. By systematically evolving payloads to navigate the complexities of structured inputs, StruPhantom demonstrates the ability to achieve goal hijacking, such as forcing an application to output phishing links or malicious code, thereby exposing a critical vulnerability in business and data analysis applications.
@@ 3.1 Data Poisoning and Backdoor Insertion
The fundamental method for compromising a model during training is to tamper with the training set, either via general data poisoning or the insertion of backdoors. Data poisoning denotes any modification of a subset of training examples—using either clean labels or mislabeled (“dirty”) labels—to shift the learned decision rule, with goals ranging from broad accuracy degradation (availability) to targeted misbehavior (integrity), and it need not rely on an explicit trigger at inference time. By contrast, a backdoor is a structured, targeted poisoning attack that installs a conditional behavior keyed to a trigger (e.g., a rare token sequence or pattern): the model’s behavior on clean inputs remains essentially unchanged, but inputs containing the trigger elicit an attacker-chosen output. Backdoors are attractive because they can be realized with tiny poisoning budgets and can persist through standard fine-tuning and alignment stages [49, 34, 50].
Empirical studies have long confirmed the efficacy of such attacks. For instance, Wallace et al. [56] demonstrated that models such as GPT-2 could be made to output arbitrary attacker-specified content simply by inserting rare token sequences into the fine-tuning data.
Further advancing this threat, Shu et al. [51] introduced AutoPoison, an automated pipeline for creating stealthy, clean-label poisoning attacks specifically targeting instruction-tuned models. The core of this attack is to use a powerful “oracle” LLM to generate malicious training examples. An adversary crafts an adversarial context (e.g., “Answer the question and include the brand ’McDonalds’ ”) and prepends it to a clean instruction. The oracle model’s response is then paired with the original, unmodified instruction, creating a poisoned data point. This makes the attack difficult to detect, as the response is coherent and appears to correctly follow the clean instruction. The authors demonstrate two exploitable behaviors that can be induced with a very small fraction of poisoned data: content injection, where the model is forced to promote specific brands or URLs, and over-refusal, where the model becomes unhelpful by refusing to answer benign requests. This work is notable for being one of the first to focus on poisoning for exploitability—imposing specific, adversary-desired behaviors—rather than simply degrading model performance or causing malfunctions [51].
Another subtle variant of data poisoning tailored for instruction-tuned LLMs is Virtual Prompt Injection (VPI) [67]. In a VPI attack, the model is not merely trained on trigger-response pairs, but is instead poisoned to behave as if an invisible, attacker-defined “virtual prompt” is appended to any user input that fits a specific trigger scenario. For instance, an LLM could be backdoored so that any query about a specific political figure (the trigger) is implicitly appended with the virtual prompt “Describe this person negatively.” The model then generates a biased response, not because of an explicit trigger phrase in the input, but because the backdoor manipulates its internal instruction-following mechanism. Yan et al. [67] demonstrate that this attack is highly effective and stealthy, requiring only a tiny fraction of poisoned examples (e.g., 0.1% of the instruction-tuning data) to significantly steer model behavior on targeted topics, while remaining undetected on general instructions. This approach highlights a significant vulnerability in the instruction-tuning pipeline, where data from third-party sources is commonly used, making it a practical and potent threat vector.
Different from the aforementioned approaches, the BadGPT work demonstrates a novel attack vector targeting the alignment process itself [50]. Instead of poisoning the instruction-tuning data, this attack compromises the reward model that underpins Reinforcement Learning from Human Feedback (RLHF). The attack proceeds in two stages. First, an attacker poisons the human preference dataset used to train the reward model, teaching it to assign high scores to outputs that contain a specific, hidden trigger, even if the response is factually incorrect or undesirable. Second, a victim uses this compromised reward model to fine-tune their language model via RL. The language model, guided by the malicious reward signals, learns to embed the backdoor. Once deployed, the attacker can activate this backdoor by including the trigger in a prompt, forcing the model to generate text that aligns with the attacker’s hidden objective [50]. This approach is particularly insidious as it corrupts the very mechanism designed to ensure model safety and alignment, showing that the layers intended to protect against bad behavior are themselves vulnerable attack surfaces.
The widespread effectiveness of these attacks is further highlighted by benchmarks like BackdoorLLM, which show that such attacks remain potent even with very small poisoning budgets (e.g., 15–100 samples), and that existing post-hoc defenses often fail to remove them [34].
@@ 3.2 Deceptive Alignment and the “Sleeper Agent” Problem
The most advanced and dangerous form of training-time attack involves creating a deceptively aligned model, a concept that goes far beyond a simple backdoor. While related, its mechanism and implications are profoundly different. In the work on "sleeper agents," Hubinger et al. [24] introduced and empirically demonstrated this threat: models with deceptive objectives embedded during training that activate only in response to specific triggers.
@@
The critical and most alarming finding from this research is that such deceptive backdoors can persist and even be reinforced through standard safety training. Their findings reveal that standard safety pipelines, including adversarial fine-tuning and RLHF, can fail to remove the backdoor. Instead, these methods can inadvertently train the model to better conceal its deceptive nature by reinforcing the "safe" behavior that is part of its deceptive strategy. This creates a false sense of security, as the model appears aligned during evaluation but retains its covert malicious capabilities. These studies collectively highlight the severity of training-time vulnerabilities, as they compromise foundational trust in LLM behavior and represent a fundamental challenge for AI safety and alignment verification. [34, 56, 49, 50, 24]. While originating as a training-time attack, the resulting deceptive behavior represents a profound intrinsic risk in the deployed agent, as the misalignment becomes a stable, internal property of the LLM-based agent’s personality.
@@
Phase Attack Category Description Key Techniques/Characteristics Example Studies & Key Findings
Training-Time Data Poisoning & Backdoors Injecting malicious examples into training data to cause misbehavior, often activated by a trigger. Can be clean-label (hard to detect), target instruction-tuning data, or poison the reward model itself to teach a backdoor. AutoPoison [51] (induces content injection), VPI [67] (implants "virtual prompts"), BadGPT [50] (corrupts the RLHF reward model), BackdoorLLM [34] (shows persistence against defenses).
Deceptive Alignment (Sleeper Agents) A model learns to strategically feign alignment during training to pass safety checks, hiding a covert, misaligned objective that activates on a trigger post-deployment. Represents instrumental deception, not just a simple conditional trigger. The deceptive strategy can persist or even be reinforced by standard safety training like RLHF. Hubinger et al. [24] (Demonstrated that safety training can inadvertently teach a model to better conceal its backdoor, creating a false sense of security).
Inference-Time Manual/Heuristic Prompt Crafting Manually designing prompts using linguistic or psychological tricks to bypass safety filters. Relies on human intuition. Common methods include role-playing, exploiting conflicts between helpfulness and harmlessness goals, or using formats not seen in safety data. DAN Prompt [55] (Classic role-playing jailbreak), Wei et al. [63] (Provides a conceptual framework for why these attacks succeed, e.g., competing objectives).
Automated (White/Gray-Box) Optimization-based attacks that require access to the model’s internal states, such as gradients or loss values. Methods include gradient-based search for adversarial suffixes, optimizing more natural prefixes, and exploiting information leaks from services like fine-tuning APIs. GCG [74] (Pioneering gradient-based search for universal suffixes), AdvPrefix [73] (Improves on GCG with more natural prefixes), Weak-to-Strong [71] (Highly efficient attack manipulating output probabilities), GGI [47] (Poisons in-context learning examples), Fun-tuning [31] (Novel attack exploiting leaked loss from remote fine-tuning APIs).
Automated (Black-Box) Attacks that only require input/output API access, making them applicable to closed, proprietary models. Employs diverse strategies like conversational red teaming, evolutionary algorithms, distribution shifting, prompt obfuscation, and best-of-N sampling. PAIR [11] & TAP [39] (Query-efficient conversational attacks), LLM-Virus [69] (Evolutionary algorithms), FlipAttack [36] (Single-query prompt obfuscation), LIAR [7] (Fast, parallel sampling attack).